OpenStudy (jenniferjuice):

Which point lies in the solution set for the following system of inequalities? y > x + 4 y y > -2x + 2

OpenStudy (mertsj):

Take each point, plug it into each inequality. Whichever point results in two true sentences is the answer.

OpenStudy (ybarrap):

add them first 2y > -x + 4y +2 2y - 4y > -x + 2 -2y > -x + 2 -y > -x/2 + 1 y < x/2 + 1 So if you have a line, y = x/2 + 1, with slope = 1/2 and y-intercept = 1, then all the Y values below this line are included in the set

OpenStudy (jenniferjuice):

(-4, 0) (-3, 2) (2, -1) (0, 6)

OpenStudy (mertsj):

Plug them in.